# thanks: Retired Professor

### v2

@@ II.2.2 The parametric microcanonical inflection point analysis
A direct analogy can be drawn between Fisher zero analysis and microcanonical methods, as mathematically evident from eqs. (5) and (24), both approaches involve investigating the behavior of a specific function as it approaches zero. Recognizing that $Z=Z(\beta)$ and ${\gamma}={\gamma}(\mathsf{e})$, i.e., they are functions of distinct variables, we propose examining both within the framework of a unified parameter.
As alluded in the introduction, we considered that eq. (18) can be solved for $\mathsf{e}$, yielding $\mathsf{e}=\mathsf{e}({\bar{\beta}})$, which allows for the derivation of parametric curves such as $\mathsf{s}=\mathsf{s}({\bar{\beta}})$, ${\gamma}={\gamma}({\bar{\beta}})$ and so forth. These parametric representations, as discussed, is not single-valued in the unstable region. Moreover, as $\bar{\beta}$ remains bounded at the transition point, eq. (24) continues to be valid for the parametric representation.
@@
Additionally, acting in accordance with earlier microcanonical analysis studies, the analysis of ${\gamma}$ indicates a loop in the parametric curve, with the knot point serving as the indicator of the transition temperature. This loop structure effectively captures the behavior associated with the transition, complementing the insights gained from the parametric curve analysis of entropy. Moreover, the loop formation is directly associated with the the undulatory curve of $\bar{\beta}(\mathsf{e})$ in the unstable region. This profile can be interpreted as a perturbation on the line of the Maxwell equal area construction for this curve. A decrease in the magnitude of this perturbation results in a concomitant reduction of the loop’s width. Consequently, in the limit of zero perturbation, the loop asymptotically approaches a vertical line (an interesting pattern that resembles the distribution of Fisher zeros in this region). It is woth mentioning that this perturbation arises from finite-size effects.
@@ III.1 Lennard-Jones Cluster
We consider $N$ particles interacting via the Lennard-Jones (LJ) potential as a case study of the first-order phase transitions. The LJ potential can be written as:
@@
where $r_{ij}=|\mathbf{r}_{j}-\mathbf{r}_{i}|$ is the distance between the particles $i$ and $j$. We chose a reduced unit system such that $\epsilon=1$ and $\sigma=2^{-1/6}$, the latter was chosen to lead the minimum of the potential at distance $r_{ij}=r_{0}=1$.
Recently, we conducted an extensive study of this model using the traditional microcanonical inflection point and the Fisher’s zeros map analysis [19], both derived from DOS obtained via the Replica Exchange Wang-Landau (REWL) method [33, 34, 35]. In this previous study, we considered $N=147$ particles confined to a sphere of radius $r_{c}=4\sigma$ to reproduce the transition temperature ($T\approx 0.36$) reported in the literature [36]. We determined the transition temperatures to be $\bar{T}_{tr}=0.3666(8)$ from the microcanonical inflection point, $\bar{T}_{tan}=0.364(1)$ from the double-tangent line construction, and $T_{1}/k_{B}=0.3622(3)$ from the leading zeros of the Fisher’s zeros map. Additionally, we demonstrated the linear behavior of the dominant zeros. It is worth mentioning that, for finite systems, different quantities provide distinct transition temperatures, which converge to a single transition value as the thermodynamic limit is approached [37]. In order to authenticate the parametric approach to the microcanonical analysis, we will reuse the same set of raw data in this present study, shown in the following Section.
To illustrate the scaling behavior of $\gamma$ described by eq. (24), we simulate $N=55$ particles inside a sphere of radius $r_{c}=3.5\sigma$, also chosen to reproduce the transition temperature ($T\approx 0.29$) reported in the literature [36] Besides that, it is claimed that the results for the LJ-cluster are independent of the volume if densities ($N/V$) are lower than that of the bulk liquid at the triple point, for the case of $147$ particles it means $r_{c}>3.7\sigma$ and for $55$ particles $r_{c}>2.6\sigma$ [38]. Therefore, to gain insight into the behavior of the Fisher’s zeros pattern along the first-order transition line, we also study a system of $N=147$ particles confined within a sphere of radius $r_{c}=3.5\sigma$. Both considerations are presented in Section III.1.2.
@@
To illustrate the proposal analysis, Fig. 1 depicts in the solid black line the entropy per spin as a function of the energy density, $\mathsf{s}(\mathsf{e})$, for the 147 particles inside a sphere of radius $r_{c}=4.0\sigma$. It is worth emphasizing that $\mathsf{s}(\mathsf{e})$ is the output of the REWL simulations. The double-tangent line construction is shown by the dotted black line. The red dashed curve in this graph represents the microcanonical inverse temperature, $\bar{\beta}(\mathsf{e})$, obtained by the derivative of the black line, as given by eq. (18). For each value of $\mathsf{e}$, we plot the ordered triple $(\mathsf{e},\bar{\beta},\mathsf{s})$, illustrated by black circles. The projection of this curve onto the $\mathsf{s}\times\bar{\beta}$-plane yields the parametric curve $\mathsf{s}(\bar{\beta}),$ shown as the dotted-dashed blue curve [^3] The projection of the triple $(\mathsf{e},\bar{\beta},\mathsf{s})$ onto the $s\times\bar{\beta}$-plane is a graphical construction. We save our data into a file where the first column represents the energy density ($\mathsf{e}$), the second column represents the entropy ($\mathsf{s}$), and the subsequent columns represent the first ($\bar{\beta}$), second ($\gamma$), and third ($\delta$) derivatives of $s$ with respect to $\mathsf{e}$. Instead of plotting $\mathsf{s}$ against $\mathsf{e}$ we plot $\mathsf{\mathsf{s}}$ against $\bar{\beta}$, where $\mathsf{\mathsf{s}}$ and $\bar{\beta}$ correspond to the same $\mathsf{e}$. This procedure is similarly applied to the other quantities. Moreover, the curve $\mathsf{s}(\mathsf{e})$ obtained from the REWL was fitted using a Bézier curve, and the derivatives of $\mathsf{s}$ were taken as the derivatives of the Bézier curve, which are also Bézier curves, see section 4.3 in the book cited in Ref. [32].. This curve is detailed in Fig. 2, where the temperature obtained from the Fisher’s zeros analysis leads to the hued regions $A_{1}\approx A_{2}$. Therefore, the temperature of the leading zero corroborates with equal area construction, which is proposed to adjust $\mathsf{s}(\bar{\beta})$ to comply with the uniqueness domain criteria of a function by eliminating the original points in the shaded regions and replace them by the vertical line. Consequently, this construction leads to the discontinuity of entropy, a defining feature in the modern classification of phase transitions. A visual inspection in Fig. 1 demonstrates that the equal area construction on $\mathsf{s}(\bar{\beta})$ is clearer than the double-tangent line construction on the convex intruder on $\mathsf{s}(\mathsf{e})$. This clarity facilitates the estimation of latent heat and aids in determining the order of the transition. Moreover, in the Fig. 2, we measured the latent heat to be $\mathcal{L}=55.1(3)$. Inserting this value into eq. (17), it leads to $\Delta\tau=\tau_{j+1}-\tau_{j}=0.1150(8)$, which differ by only $4.5\%$ from the average of the distances between the dominant zeros, $\langle\Delta\tau\rangle=0.110(2)$, measured on the zeros maps [19]. This result further supports the proposed approach.
@@
Figure 2: (Color online) Parametric curve defined by the entropy per particles, $\mathsf{s}(\mathsf{e})$, and the microcanonical inverse temperature, $\bar{\beta}(\mathsf{e})$, for the 147-LJ cluster. The inset provides a broader view of the curve. The error bars are smaller than the symbol, then not shown. The hued regions $A_{1}\approx A_{2}$ represent the equal areas construction. The vertical lines indicate estimations of the transition temperature: $\bar{\beta}_{tr}=2.728(6)$, shown as the double-dotted-dashed magenta line, obtained from the regular microcanonical inflection point; $\bar{\beta}_{tan}=2.751(9)$, shown as the dashed blue line, determined by the double-tangent line construction; and $k_{B}\beta_{1}=2.761(2)$, shown as the dotted-dashed green line, identified by the Fisher´s zeros analysis. These estimations are from Ref [19]. The inverse temperature at the knot position is represented by the solid red line, $\beta_{knot}=2.732(1)$.
Fig. 3 demonstrates a similar process for obtaining the parametric curve $\gamma(\bar{\beta})$, which is shown in detail in Fig. 4. Additionally, Fig. 3 illustrates the regular microcanonical inflection point analysis by the dotted black line. This line projects the energy of the peak position of $\gamma(\mathsf{e})$ onto $\bar{\beta}(\mathsf{e})$, leading to $\bar{\beta}_{tr}=\bar{\beta}(\mathsf{e}_{tr})$. The loop on the parametric curve $\gamma(\bar{\beta})$ for the first-order transition is illustrated in this figure. By applying the uniqueness domain criterion for functions, the curve is truncated at the knot position, thereby eliminating the loop points and defining the transition temperature at this position. In Fig. 4 we measured the temperature of the knot position to be $\bar{T}_{knot}=0.3660(1)$. It is worth mentioning that $|\bar{T}_{tr}-\bar{T}_{knot}|$ is smaller than the error of $\bar{T}_{tr}$, where $\bar{T}_{tr}$ is the transition temperature obtained from the regular microcanonical analysis. Additionally, the latent heat was calculated using all estimated transition temperatures, with the results differing within the error bars. Specifically, in addition to the previously mentioned $\mathcal{L}_{1}=55.1(3)$ obtained from $\beta_{1}$, we find that $\beta_{tr}$ yields $\mathcal{L}_{tr}=55.0(6)$, $\beta_{tan}$ estimates $\mathcal{L}_{tan}=54.2(7)$, and $\beta_{knot}$ measures $\mathcal{L}_{knot}=54.7(2)$. This yields an average value of $\mathcal{L}_{avg}=54.8(5)$.
@@
Fig. 5 depicts the scaling behavior of $\gamma(\bar{\beta})$ with system size. The data for $L=55$ particles confined within a spherical volume of radius $r_{c}=3.5\sigma$ are represented by black circles. The knot position corresponds to the transition temperature, $\bar{T}_{knot}=0.2982(2)$, as indicated by the solid red line. For comparison, the transition temperature associated with dominant zeros is shown by the dashed magenta line. Results for $N=147$ particles are displayed as red squares for $r_{c}=3.5\sigma$ and green diamonds for $r_{c}=4.0\sigma$.
A clear shrinkage of the loop towards $\gamma=0$ is observed with increasing system size, consistent with the expectation from eq. (24). Notably, the loop is slightly smaller for $L=147$ at $r_{c}=4.0\sigma$ compared to $r_{c}=3.5\sigma$, as evident in the inset of the figure. It is well-established that for simple systems, the first-order transition line tends towards a critical point with increasing pressure and, consequently, a decrease in latent heat. A systematic investigation of this aspect falls outside the scope of the present study. Primary results on highly compressed configurations have exhibited complex behavior characterized by two distinct transition signals, resembling observations for LJ clusters with an intermediate number of particles between the established “magic numbers” (the chosen values of $N=55$ and $N=147$ are recognized as magic numbers). While LJ clusters corresponding to magic numbers typically display a strong first-order transition from the liquid phase to the energetically preferred low-temperature Mackay icosahedral solid package, intermediate-sized clusters display an additional solid-solid transition. This interjacent transition occurs between anti-Mackay (hexagonal close-packed, HCP, overlayers) and Mackay (face-centered cubic, FCC, overlayers) icosahedral shell packaging [40].
@@
Fig. 6 presents the Fisher’s zeros map for the Lennard-Jones (LJ) cluster constrained to a sphere of radius $r_{c}=3.5\sigma$. We use MPSolve [41, 42] as the root finder. A characteristic pattern of equally spaced vertical lines of zeros is observed, converging towards the real temperature axis, indicative of a first-order phase transition. Panel (a) displays the map for $N=55$ particles, where the real part of the leading zero corresponds to a transition temperature of $k_{B}T_{1}=0.2956(4)$ and latent heat of $\mathcal{L}=30.80(3)$. Panel (b) shows the map for $N=147$ particles, with a measured transition temperature of $k_{B}T_{1}=0,3676(4)$ and latent heat of $\mathcal{L}=51.51(3)$. The zeros map for a 147-particle LJ cluster confined within a sphere of radius $r_{c}=4.0\sigma$ can be found in reference [19]. Each symbol in the figure represents a map obtained from an independent simulation, it can be noted that, as it is well-known, the zeros of the partition function are highly sensitive to statistical fluctuations, except for the dominant zeros.
@@
The two-dimensional (2D) Lenz-Ising model serves as a theoretical framework for the second-order phase transitions [43]. This model considers spin-$1/2$ particles, arranged in a fixed lattice, that interacts with their nearest neighbors. Lars Onsager derived the exact solution for the square lattice in 1944 [44], where the critical temperature is deduced to be $T_{c}=2/\ln(1+\sqrt{2})$.

### v4

@@ II.2.2 The parametric microcanonical inflection point analysis
A direct analogy can be drawn between Fisher zero analysis and microcanonical methods, as mathematically evident from Eqs. (5) and (28); both approaches involve investigating the behavior of a specific function as it approaches zero. Recognizing that $Z=Z(\beta)$ and ${\gamma}={\gamma}(\mathsf{e})$, i.e., they are functions of distinct variables, we propose examining both within the framework of a unified parameter.
As alluded to in the introduction, we considered that Eq. (18) can be solved for $\mathsf{e}$, yielding $\mathsf{e}=\mathsf{e}({\bar{\beta}})$, which allows for the derivation of parametric curves such as $\mathsf{s}=\mathsf{s}({\bar{\beta}})$, ${\gamma}={\gamma}({\bar{\beta}})$ and so forth. These parametric representations, as discussed, are not single-valued in the unstable region. Moreover, as $\bar{\beta}$ remains bounded at the transition point, Eq. (28) continues to be valid for the parametric representation.
@@
Additionally, acting in accordance with earlier microcanonical analysis studies, the analysis of ${\gamma}$ indicates a loop in the parametric curve, with the knot point serving as the indicator of the transition temperature. This loop structure effectively captures the behavior associated with the transition, complementing the insights gained from the parametric curve analysis of entropy. Moreover, the loop formation is directly associated with the undulatory curve of $\bar{\beta}(\mathsf{e})$ in the unstable region. This profile can be interpreted as a perturbation on the line of the Maxwell equal area construction for this curve. A decrease in the magnitude of this perturbation results in a concomitant reduction of the loop’s width. Consequently, in the limit of zero perturbation, the loop asymptotically approaches a vertical line (an interesting pattern that resembles the distribution of Fisher zeros in this region). It is worth mentioning that this perturbation arises from finite-size effects.
@@ III.1 Lennard-Jones Cluster
We consider $N$ particles interacting via the Lennard-Jones (LJ) potential as a case study of the first-order phase transitions. The LJ potential can be written as
@@
where $r_{ij}=|\mathbf{r}_{j}-\mathbf{r}_{i}|$ is the distance between the particles $i$ and $j$. We chose a reduced unit system such that $\epsilon=1$ and $\sigma=2^{-1/6}$; the latter was chosen to lead the minimum of the potential at distance $r_{ij}=r_{0}=1$.
Recently, we conducted an extensive study of this model using the traditional microcanonical inflection point and the Fisher’s zeros map analysis [21], both derived from the DOS obtained via the replica exchange Wang-Landau method [39, 40, 41]. In this previous study, we considered $N=147$ particles confined to a sphere of radius $r_{c}=4\sigma$ to reproduce the transition temperature ($T\approx 0.36$) reported in the literature [42]. We determined the transition temperatures to be $\bar{T}_{tr}=0.3666(8)$ from the microcanonical inflection point, $\bar{T}_{tan}=0.364(1)$ from the double-tangent line construction, and $T_{1}/k_{B}=0.3622(3)$ from the leading zeros of the Fisher’s zeros map. Additionally, we demonstrated the linear behavior of the dominant zeros. It is worth mentioning that, for finite systems, different quantities provide distinct transition temperatures, which converge to a single transition value as the thermodynamic limit is approached [43]. In order to authenticate the parametric approach to the microcanonical analysis, we will reuse the same set of raw data in this present study, shown in the following Section.
To illustrate the scaling behavior of $\gamma$ described by Eq. (28), we simulate $N=55$ particles inside a sphere of radius $r_{c}=3.5\sigma$, also chosen to reproduce the transition temperature ($T\approx 0.29$) reported in the literature [42] Besides that, it is claimed that the results for the LJ-cluster are independent of the volume if densities ($N/V$) are lower than that of the bulk liquid at the triple point, for the case of $147$ particles it means $r_{c}>3.7\sigma$ and for $55$ particles $r_{c}>2.6\sigma$ [44]. Therefore, to gain insight into the behavior of the Fisher’s zeros pattern along the first-order transition line, we also study a system of $N=147$ particles confined within a sphere of radius $r_{c}=3.5\sigma$. Both considerations are presented in Sec. III.1.2.
@@
To illustrate the proposal analysis, Fig. 1 depicts in the solid black line the entropy per spin as a function of the energy density, $\mathsf{s}(\mathsf{e})$, for the 147 particles inside a sphere of radius $r_{c}=4.0\sigma$. It is worth emphasizing that $\mathsf{s}(\mathsf{e})$ is the output of the REWL simulations  [^3] The data are publicly available on the zenodo.org[77]. The data presented here are the mean values of five independent simulations. Error bars were estimated via the jackknife resampling method [46], and they are not shown when smaller than the symbol size. The double-tangent line construction is shown by the dotted black line. The red dashed curve in this graph represents the microcanonical inverse temperature, $\bar{\beta}(\mathsf{e})$, obtained by the derivative of the black line, as given by Eq. (18). For each value of $\mathsf{e}$, we plot the ordered triple $(\mathsf{e},\bar{\beta},\mathsf{s})$, illustrated by black circles. The projection of this curve onto the $\mathsf{s}\times\bar{\beta}$-plane yields the parametric curve $\mathsf{s}(\bar{\beta}),$ shown as the dotted-dashed blue curve  [^4] The projection of the triple $(\mathsf{e},\bar{\beta},\mathsf{s})$ onto the $s\times\bar{\beta}$-plane is a graphical construction. We save our data into a file where the first column represents the energy density ($\mathsf{e}$), the second column represents the entropy ($\mathsf{s}$), and the subsequent columns represent the first ($\bar{\beta}$), second ($\gamma$), and third ($\delta$) derivatives of $s$ with respect to $\mathsf{e}$. Instead of plotting $\mathsf{s}$ against $\mathsf{e}$ we plot $\mathsf{\mathsf{s}}$ against $\bar{\beta}$, where $\mathsf{\mathsf{s}}$ and $\bar{\beta}$ correspond to the same $\mathsf{e}$. This procedure is similarly applied to the other quantities. Moreover, the curve $\mathsf{s}(\mathsf{e})$ obtained from the REWL was fitted using a Bézier curve, and the derivatives of $\mathsf{s}$ were taken as the derivatives of the Bézier curve, which are also Bézier curves, see section 4.3 in the book cited in Ref. [38].. This curve is detailed in Fig. 2, where the temperature obtained from the Fisher’s zeros analysis leads to the hued regions $A_{1}\approx A_{2}$. Therefore, the temperature of the leading zero corroborates with equal area construction, which is proposed to adjust $\mathsf{s}(\bar{\beta})$ to comply with the uniqueness domain criteria of a function by eliminating the original points in the shaded regions and replacing them with the vertical line. Consequently, this construction leads to the discontinuity of entropy, a defining feature in the modern classification of phase transitions. A visual inspection in Fig. 1 demonstrates that the equal area construction on $\mathsf{s}(\bar{\beta})$ is clearer than the double-tangent line construction on the convex intruder on $\mathsf{s}(\mathsf{e})$. This clarity facilitates the estimation of latent heat and aids in determining the order of the transition. Moreover, in Fig. 2, we measured the latent heat to be $\mathcal{L}=55.1(3)$. Inserting this value into Eq. (17), leads to $\Delta\tau=\tau_{j+1}-\tau_{j}=0.1150(8)$, which differ by only $4.5\%$ from the average of the distances between the dominant zeros, $\langle\Delta\tau\rangle=0.110(2)$, measured on the zeros maps [21]. This result further supports the proposed approach.
@@
Figure 2: (Color online) Parametric curve defined by the entropy per particles, $\mathsf{s}(\mathsf{e})$, and the microcanonical inverse temperature, $\bar{\beta}(\mathsf{e})$, for the 147-LJ cluster. The inset provides a broader view of the curve. The error bars are smaller than the symbol, then not shown. The hued regions $A_{1}\approx A_{2}$ represent the equal areas construction. The vertical lines indicate estimations of the transition temperature: $\bar{\beta}_{tr}=2.728(6)$, shown as the double-dotted-dashed magenta line, obtained from the regular microcanonical inflection point; $\bar{\beta}_{tan}=2.751(9)$, shown as the dashed blue line, determined by the double-tangent line construction; and $k_{B}\beta_{1}=2.761(2)$, shown as the dotted-dashed green line, identified by the Fisher´s zeros analysis. These estimations are from Ref. [21]. The inverse temperature at the knot position is represented by the solid red line, $\beta_{knot}=2.732(1)$.
Figure 3 demonstrates a similar process for obtaining the parametric curve $\gamma(\bar{\beta})$, which is shown in detail in Fig. 4. Additionally, Fig. 3 illustrates the regular microcanonical inflection point analysis by the dotted black line. This line projects the energy of the peak position of $\gamma(\mathsf{e})$ onto $\bar{\beta}(\mathsf{e})$, leading to $\bar{\beta}_{tr}=\bar{\beta}(\mathsf{e}_{tr})$. The loop on the parametric curve $\gamma(\bar{\beta})$ for the first-order transition is illustrated in this figure. By applying the uniqueness domain criterion for functions, the curve is truncated at the knot position, thereby eliminating the loop points and defining the transition temperature at this position. In Fig. 4, we measured the temperature of the knot position to be $\bar{T}_{knot}=0.3660(1)$. It is worth mentioning that $|\bar{T}_{tr}-\bar{T}_{knot}|$ is smaller than the error of $\bar{T}_{tr}$, where $\bar{T}_{tr}$ is the transition temperature obtained from the regular microcanonical analysis. Additionally, the latent heat was calculated using all estimated transition temperatures, with the results differing within the error bars. Specifically, in addition to the previously mentioned $\mathcal{L}_{1}=55.1(3)$ obtained from $\beta_{1}$, we find that $\beta_{tr}$ yields $\mathcal{L}_{tr}=55.0(6)$, $\beta_{tan}$ estimates $\mathcal{L}_{tan}=54.2(7)$, and $\beta_{knot}$ measures $\mathcal{L}_{knot}=54.7(2)$. This yields an average value of $\mathcal{L}_{avg}=54.8(5)$.
@@
Figure 5 depicts the scaling behavior of $\gamma(\bar{\beta})$ with system size. The data for $L=55$ particles confined within a spherical volume of radius $r_{c}=3.5\sigma$ are represented by black circles. The knot position corresponds to the transition temperature, $\bar{T}_{knot}=0.2982(2)$, as indicated by the solid red line. For comparison, the transition temperature associated with dominant zeros is shown by the dashed magenta line. Results for $N=147$ particles are displayed as red squares for $r_{c}=3.5\sigma$ and green diamonds for $r_{c}=4.0\sigma$.
A clear shrinkage of the loop toward $\gamma=0$ is observed with increasing system size, consistent with the expectation from Eq. (28). Notably, the loop is slightly smaller for $L=147$ at $r_{c}=4.0\sigma$ compared to $r_{c}=3.5\sigma$, as evident in the inset of the figure. It is well established that for simple systems, the first-order transition line tends toward a critical point with increasing pressure and, consequently, a decrease in latent heat. A systematic investigation of this aspect falls outside the scope of the present study. Primary results on highly compressed configurations have exhibited complex behavior characterized by two distinct transition signals, resembling observations for LJ clusters with an intermediate number of particles between the established “magic numbers” (the chosen values of $N=55$ and $N=147$ are recognized as magic numbers). While LJ clusters corresponding to magic numbers typically display a strong first-order transition from the liquid phase to the energetically preferred low-temperature Mackay icosahedral solid package, intermediate-sized clusters display an additional solid-solid transition. This interjacent transition occurs between anti-Mackay (hexagonal close-packed, HCP, overlayers) and Mackay (face-centered cubic, FCC, overlayers) icosahedral shell packaging [48].
@@
Figure 6 presents the Fisher’s zeros map for the Lennard-Jones (LJ) cluster constrained to a sphere of radius $r_{c}=3.5\sigma$. We use MPSolve [49, 50] as the root finder. A characteristic pattern of equally spaced vertical lines of zeros is observed, converging toward the real temperature axis, indicative of a first-order phase transition. Panel (a) displays the map for $N=55$ particles, where the real part of the leading zero corresponds to a transition temperature of $k_{B}T_{1}=0.2956(4)$ and latent heat of $\mathcal{L}=30.80(3)$. Panel (b) shows the map for $N=147$ particles, with a measured transition temperature of $k_{B}T_{1}=0,3676(4)$ and latent heat of $\mathcal{L}=51.51(3)$. The zeros map for a 147-particle LJ cluster confined within a sphere of radius $r_{c}=4.0\sigma$ can be found in reference [21]. Each symbol in the figure represents a map obtained from an independent simulation. It can be noted that, as it is well-known, the zeros of the partition function are highly sensitive to statistical fluctuations, except for the dominant zeros.
@@
The two-dimensional (2D) Lenz-Ising model serves as a theoretical framework for the second-order phase transitions [51]. This model considers spin-$1/2$ particles, arranged in a fixed lattice, that interact with their nearest neighbors. Lars Onsager derived the exact solution for the square lattice in 1944 [6], where the critical temperature is deduced to be $T_{c}=2/\ln(1+\sqrt{2})$.